# Exploring the Zero-Shot Capabilities of LLMs Handling Multiple Problems at once

### v2

@@ 3.3 Experimental Setups
3.4 Results
Overall, we observe that although LLMs are competent zero-shot multi-problem solvers for classification (Section 3.4.1), their performance decreases significantly when the same sets of multiple problems are presented differently (Section 3.4.2).
3.4.1 Zero-shot MPP: SingleClf vs. BatchClf
In line with previous studies Cheng et al. (2023); Lin et al. (2024) on few-shot MPP, we find that zero-shot MPP also performs comparably well or even better than zero-shot SPP in a more cost-efficient way. We further validate this finding by comparing the errors LLMs make under SPP and MPP.
Although the BatchClf accuracy generally declines as the task size increases (Fig 2), all 7 LLMs achieve accuracy of at least 90% that of SingleClf across the benchmarks most of the time (see Table 5 in Appendix B). Overall, the SingleClf accuracy for the 7 LLMs on the 6 benchmarks is 75.5% and the BatchClf accuracy is 72.3%, a minor 3.2% absolute drop from the former.
Zero-shot MPP can be cost-efficient.
With a general task instruction being shared, zero-shot MPP avoids repeating this instruction for each problem to be solved. Given the strong zero-shot MPP performance, this makes MPP a promising cost-efficient zero-shot prompting technique. To illustrate, we choose for each model/task combination the largest BatchClf task size that achieves at least 95% SingleClf accuracy for that pair. We observe that MPP reduces substantial inference costs for all LLMs run on the 6 benchmarks, ranging from from 30.7% to 82.0% (see Fig 7 in Appendix B).
LLMs make similar errors evenly across positions under multiple problems.
Given the results above, two natural questions arise: do LLMs make similar errors under MPP and how do their errors distribute across positions? For each one of the 180 LLM/benchmark/task size combinations, we use chi-squared tests to compare the proportional error distribution across class labels under SingeClf and BatchClf and compute the cumulative error density across positions under BatchClf. We observe that (1) only in 9 out of 180 (or 5%) cases, error labels are distributed significantly differently between SingleClf and BatchClf (p < 0.05); and (2) surprisingly, LLMs typically do not display a clear positional bias or a serial position effect as known in psychology Murdock (1962), when solving sufficiently many problems at once (Fig 3). This is in contrast to previous studies based on single-problem prompts where LLMs are found to be better at using information from the beginning (primacy bias) or the end (receny bias) of the prompt Liu et al. (2024); Levy et al. (2024).
We hypothesize that the differences in observations may be due to different ways of measuring positional biases. Without a precise definition of positional biases and clear notions of what counts as “beginning,” “middle,” and “end,” we find it necessary to look at cumulative error density (Fig 3) to identify the overall positional error distribution. We leave the arguments for future studies.
3.4.2 MPP Reformulations: Selection Tasks
We further compare LLM’s performance between BatchClf and the two selection tasks.
@@ 3.4.2 MPP Reformulations: Selection Tasks
In our experiments, LLMs nearly always perform much better in BatchClf than in SelectOne and SelectAll under the same conditions with a consistent and stable margin, even when the task size is just 3 or 5 (Figs 12). The overall discrepancy in accuracy between BatchClf and the two tasks is large and statistically significant (32% for SelectOne and 10% for SelectAll, see Table 2) and generally increases with a larger task size (Fig 2).
The sharp drop in accuracy may not be human-like, because arguably, humans should at least be able to classify and select a small number (e.g., 3/5) of texts equally well simply by thinking over the problems (i.e., zero-shot). This suggests either LLMs’ lack of true understanding of the problems presented in various forms or their struggle at index selection.
LLMs lack internalized planning and consistently select inconsistently.
@@
We investigate the reasons in Fig 4, which shows that when asked to select text indices for one class label at a time independently, LLMs almost always assign an index to multiple labels (i.e., contradiction) and leave some indices unselected (i.e., non-excluded middles) more often. This showcases a lack of internalized planning with modern zero-shot LLMs, although different LLMs may make these two types of errors in different proportions. In contrast, when LLMs have to select indices for all labels at once, they are less likely to generate directly illogical answers in a single output as their answer to the $(i+1)_{th}$ label is conditioned by their answer to the $i_{th}$ label. This highlights the purely probabilistic aspect of LLM answer generation.
LLMs can do index selection, but fail to combine it with text classification adequately.
To see if LLMs’ significantly poorer performance with selection tasks is due to their struggle at index selection, we perform a follow-up study. More concretely, we simplify the two index selection tasks by directly replacing each line of text with its gold class label in the prompts for AGNews, which has most labels. We then ask the LLMs to select indices of lines containing each label with minimal modifications to the original task instructions (see Appendix E).
Figure 5: Average accuracy of the 7 LLMs on the two simplified index selection tasks based on AGNews. We provide the original results from BatchClf, SelectOne, and SelectAll for easy comparisons.
Fig 5 shows that LLMs can perform in the simplified selection tasks much better than the original ones, with Llama-3 70B, GPT-3.5, and GPT-4 even achieving (nearly) 100% accuracy across task sizes. Their overall performance in the simplified selection tasks even surpasses their overall BatchClf accuracy. The previously observed performance gap between SelectOne and SelectAll also nearly disappears (see Appendix B for full results). In general, LLMs can indeed do index selection.
In conclusion, it appears that the tested LLMs perform less well on SelectOne and SelectAll as compared to BatchClf because they cannot adequately combine the index selection task and the classification tasks in response to a zero-shot prompt. [^1] We also tried 1-shot-CoT to explicitly prompt LLMs to do BatchClf first and then select indices from there. Although that helps, this basic finding remains. See Appendix B.6. Put differently, the LLMs lack true understanding of the problems presented in different forms, even when the number of problems is quite small (e.g., 3 or 5), which may not be human-like.
4 MPP: Reasoning
We provide the MPP experimental details and results for reasoning problems in this section.
4.1 Data
We further experiment with another 12 reasoning benchmarks (listed and referenced in Table 1), commonly used in LLM evaluation (Kojima et al., 2022; Wei et al., 2023; Zhang et al., 2023, inter alia). These benchmarks test symbolic reasoning (Coin Flips & Last Letters), commonsense reasoning (StrategyQA, CommonsenseQA, Object tracking, & Bigbench date), and arithmetic reasoning (AQuA, SVAMP, GSM8K, MultiArith, AddSub, & SingleEq), and require three answer formats (Yes/No, multiple choice, and free-response).
4.2 Evaluation Tasks
We first test on all the reasoning problems in each benchmark to establish LLM SPP baselines. Two MPP evaluation task types are designed, i.e., single-source and mixed-source multi-problem reasoning, or MultiReasonSS and MultiReasonMS. For both task types, we simply bundle multiple reasoning problems together with an indexed prefix “Qi: ,” where i is the index of each problem starting from “Q1.” We use two headers, namely, “Questions,” followed by the bundled questions, and “Answers,” followed by the answers to be generated by LLMs.
Besides vanilla zero-shot prompting, we also perform zero-shot-CoT following Kojima et al. (2022). [^2] In our early experiments, we found that zero-shot-CoT did not lead to different responses for the classification-related MPP tasks probably due to their novelty, so it was not used. Inspired by Cheng et al. (2023) as well as to control for the number of prompts generated, we consider smaller task sizes from 2 to 10. To ensure a reliable evaluation (e.g., sufficient parsable outputs), we increase the number of prompts from 100 to 300 for each benchmark/task size combination. This results in 21,600 and 18,000 prompts for MultiReasonSS and MultiReasonMS, respectively.
MultiReasonSS
For each one of the 12 benchmarks, we first sample 300 distinct sets of 10 problems. We then use the first 2, 5, and 10 problems from each problem set to make prompts.
MultiReasonMS
To robustly examine an LLM’s performance on mixed-source prompts, we create 6 distinct benchmark combinations based on the 12 benchmarks, each consisting of 10 different benchmarks. For each combination, we randomly sample 300 distinct sets of 10 problems, each coming from one benchmark and with the order of benchmarks randomized across the 300 problem sets. We then take out 2/4/6/8/10 problems from the first 2/4/6/8/10 benchmarks for each benchmark combination across each problem set to make mixed-source prompts. See Appendix B for details.
4.3 Experimental Setups
LLMs and evaluation settings
For budget reasons, we only use two LLMs, i.e., GPT-3.5 and Llama-3 70B. Since Llama-3 models tend to produce reasoning steps even when not instructed to do so, we only prompt GPT-3.5 with zero-shot-CoT. We always prompt with greedy decoding.
Performance metric
We use the per-problem accuracy (PPA), defined in Equation 1, to measure the model performance. For mixed-source prompts containing $k$ problems evenly sampled from $k$ benchmarks, we compute the expected PPA by averaging over the observed SPP performance for each one of these benchmarks.
4.4 Results
@@
We observe that although LLMs can be competent zero-shot multi-problem solvers for reasoning, their performance becomes consistently worse than expected under multiple mixed-source problems. Similar to our arguments in Section 3, the consistent performance declines even with a small number (e.g., 2 or 4) of mixed-source problems may indicate a lack of human-like understanding, as LLMs’ reasoning capabilities are easily impacted by the mixing of problems from different sources.
@@
Figure 6: Average accuracy of GPT-3.5 and Llama-3 70B on multiple single-source (MultiReasonSS) and mixed-source (MultiReasonMS) reasoning problems. (A) MultiReasonSS results on 6 selected benchmarks for space reasons. The other 6 benchmarks show similar results (see Appendix B). (B) MultiReasonMS results across 6 distinct benchmark combinations, where each benchmark contributes a problem to each mixed-source prompt.
@@
5 Exploring Model-level Factors that may Enable MPP
Since so far we have only tested decoder-only and instruction-tuned LLMs, which all show strong performance under MPP, we explore if these two factors enable MPP. For these reasons, we test with greedy decoding (1) three FLAN-T5 models Chung et al. (2022), i.e., Large, XL, XXL; and (2) three pretrained LLMs, i.e., Llama-3 8B (Base, Meta, 2024), GPT-3 1.3B, and GPT-3 175B Brown et al. (2020). We run these 6 LLMs on CoLA at task sizes 1 and 5 under zero-shot settings with results shown in Table 3. Experiments with Coin Flips show similar results, but the LLM outputs are less meaningful, as discussed in Appendix D.
We make three observations from Table 3. First, instruction tuning helps, since pretraned decoder-only LLMs either cannot handle multiple problems at once or their performance is much worse than their instruction-tuned counterparts. However, unlike what Laskar et al. (2023) suggests, instruction tuning may not be a necessary condition for MPP, as both Llama-3 8B and GPT-3 175B can perform reasonably well in BatchClf on CoLA.
Second, FLAN-T5 can barely respond to MPP, regardless of model sizes. We suspect that this is may not be due to their encoder-decoder structures, but other factors such as training data and reinforcement learning from human feedback Christiano et al. (2017). We leave it for future investigation.
Third, scaling model size seems helpful. With other factors being identical, larger models appear to perform better than the smaller ones under MPP.
@@ 5 Exploring Model-level Factors that may Enable MPP
Table 3: SingleClf and BatchClf (task size 5) accuracy (%) of three pretrained LLMs and three FLAN-T5 models on CoLA. We also include the results of Llama-3 8B (Instruct) and GPT-3.5 (likely base model: GPT-3 175B) from Section 3 to compare with their respective base models. The last column is the average number of LLM-generated answers for BatchClf (expects 5). When it is 1, accuracy is not calculated to avoid overinterpretation.
@@ 6 Conclusion
In this study, we present a comprehensive evaluation of LLMs with zero-shot MPP. We evaluate various LLMs from 4 model families on single-source multi-problem prompts constructed from 6 classification and 12 reasoning benchmarks. In line with previous few-shot results, we confirm that LLMs are competent multi-problem solvers for classification and reasoning under zero-shot settings. Moreover, we find multiple pieces of evidence that validate the strong innate multiple problem handling capabilities of LLMs, such as the similar classification errors LLMs make under SPP and MPP, the lack of obvious positional biases, and the transferrability of zero-shot-CoT under MPP.
@@
As a result of our study, we create a new benchmark comprising 53,100 zero-shot multi-problem prompts. We call it ZeMPEB, which stands for Zero-shot Multi-Problem Evaluation Benchmark. We will release ZeMPEB to aid future research.
@@ Appendix A LLM Details
We use a total of 13 LLMs in our study, from Section 3 to Section 5. Table 4 describes the specific versions for these LLMs and highlights their differences in terms of architecture, open weights, Supervised Fine-Tuning (SFT, Wei et al., 2022), and Reinforment Learning from Human Feedback (RLHF, Christiano et al., 2017) etc.
@@
Appendix B MPP: Classification
This section contains additional details for Section 3. The prompt details for the experiments can be found in Appendix E.
@@ B.1 Full Results
The full results obtained from Section 3 are visualized in Fig 8. We exclude the results of Vicuna on AGNews when task size is 100 because the prompts exceed the model’s context length.
@@ B.3 Zero-shot MPP can be cost-efficient
B.4 Positional Errors under BatchClf
B.5 Simplified Index Selection Tasks
Fig 10 shows the full results for the two simplified index selection tasks, along with results for the two original selection tasks and BatchClf, based on AGNews. Clearly, (1) all LLMs perform much better in the simplified tasks than the original ones; (2) Mixtral 8X7B, Llama-3 70B, GPT-3.5, and GPT-4 can do these simplified selection tasks even better than their BatchClf performance consistently, with the latter three achieving nearly 100% accuracy across task sizes; (3) the task complexity hierarchy between SelectOne and SelectAll nearly disappears for all LLMs except Vicua, implying that it may be challenging for Vicuna to perform index selection tasks in general.
@@ B.5 Simplified Index Selection Tasks
Figure 10: Full results for the two simplified index selection tasks versus the original tasks based on AGNews.
@@
Task SelectOne + CoT SelectAll + CoT BatchClf
Model
Vicuna 23.0 25.8 53.2 57.6 68.7
Mistral 7B 38.3 47.5 56.8 60.5 71.2
Mixtral 8x7B 47.4 38.7 65.1 58.4 73.4
Llama-3 8B 39.0 41.8 62.4 59.0 73.3
Llama-3 70B 59.4 67.1 72.9 79.2 79.4
GPT-3.5 45.5 47.5 66.7 66.3 71.9
GPT-4 66.3 71.8 78.8 81.8 81.9
Overall 45.5 48.6 65.1 66.1 74.3
Table 6: Aggregate average accuracy of SelectOne and SelectAll with and without 1-shot-CoT for each LLM. BatchClf performance is also provided for comparisons.
B.6 CoT Helps the Index Selection Tasks Slightly but the Gap Remains
We use 1-shot-CoT to prompt LLMs to do BatchClf first and then perform the two index selection tasks on the 6 benchmarks with a fixed task size 10. Table 6 shows that while overall LLMs benefit from CoT for both SelectOne and SelectAll, the benefits are generally larger for SelectOne than SelectAll (3.1% versus 1.0% improvement) and not consistent across models with Mixtral 8x7B performing even worse with CoT. The full results provided in Table D also show mixed effects across benchmarks. Interestingly, however, the task complexity hierarchy among BatchClf, SelectOne, and SelectAll remains in virtually all cases. We leave to future studies the investigation of how CoT may further help with the two selection tasks.
Appendix C MPP: Reasoning
This section provides additional details for Section 4. The prompt details for the experiments can be found in Appendix E.
@@ C.2 More Single-Source Results
Appendix D Additional Results for Section 5
This section describes the results of the three pretrained base LLMs and three FLAN-T5 models on Coin Flips at task sizes 1 and 2 from Section 5, shown in Table 7. Similar to Table 3, Table 7 shows that instruction-tuned models perform much better on multi-problem prompts and that FLAN-T5 models can barely handle multi-problem prompts.
@@ Appendix D Additional Results for Section 5
Benchmark SST-2 CoLA AGNews MRPC SNLI WiC
Model Task
Vicuna SelectOne 38.5 22.2 16.8 15.8 9.9 34.5
+ CoT 52.6 53.9 12.1 12.1 11.5 12.5
SelectAll 66.7 65.7 42.2 67.4 28.2 48.8
+ CoT 81.7 69.2 40.9 63.9 40.9 48.9
BatchClf 89.1 73.2 70.4 66.7 59.3 53.4
[…]
+ CoT 66.5 55.1 44.8 59.4 29.5 29.5
SelectAll 68.3 63.6 48.2 58.4 54.7 47.9
+ CoT 74.4 68.2 57.0 68.5 47.6 47.3
BatchClf 79.6 76.8 78.1 74.0 69.1 49.6
[…]
+ CoT 41.5 67.3 12.9 48.4 25.1 37.2
SelectAll 85.3 68.4 62.9 65.9 56.2 52.0
+ CoT 49.0 67.4 57.8 66.5 58.9 51.1
BatchClf 80.2 80.9 71.6 70.8 73.3 63.7
[…]
+ CoT 85.0 48.9 22.1 50.7 22.3 22.0
SelectAll 85.0 62.9 59.6 64.8 50.1 52.2
+ CoT 85.8 66.5 45.3 64.5 39.8 52.3
BatchClf 92.0 76.8 79.5 62.8 69.2 59.5
[…]
+ CoT 94.7 79.7 75.2 67.9 69.4 15.8
SelectAll 87.3 79.6 76.2 71.7 68.6 53.9
+ CoT 96.1 83.7 85.5 73.9 77.3 58.4
BatchClf 96.4 82.8 86.8 73.5 75.2 61.7
[…]
+ CoT 80.2 71.7 57.9 26.8 26.5 22.1
SelectAll 83.1 79.7 68.3 61.5 55.8 51.6
+ CoT 89.5 77.5 58.8 70.2 50.5 51.2
BatchClf 95.0 81.8 86.4 52.7 57.4 58.0
[…]
+ CoT 84.5 81.9 67.8 60.1 77.0 59.6
SelectAll 92.4 84.0 79.2 73.5 80.8 62.8
+ CoT 91.3 83.7 85.4 76.2 83.3 71.0
BatchClf 96.0 84.0 86.7 72.6 83.8 68.4
Overall SelectOne 68.6 58.0 48.7 37.7 27.6 32.7
+ CoT 72.1 65.5 41.8 46.5 37.3 28.4
SelectAll 81.2 72.0 62.4 66.2 56.3 52.7
+ CoT 81.1 73.7 61.5 69.1 56.9 54.3
BatchClf 89.8 79.5 79.9 67.6 69.6 59.2
Table 8: Full 1-shot-CoT results for all benchmarks and LLMS. The task size is fixed at 10.

### v3

@@ 4.3 Classification-Related Results
In our experiments, LLMs nearly always perform much better in BatchClf than in SelectOne and SelectAll under the same conditions with a consistent and stable margin, even when the task size is just 3 or 5 (Fig 1). The overall discrepancy in accuracy between BatchClf and the two tasks is large and statistically significant (32% for SelectOne and 10% for SelectAll, see Table 2) and generally increases with a larger task size (Fig 1).
The sharp drop in accuracy may not be human-like, because arguably, humans should at least be able to classify and select a small number (e.g., 3/5) of texts equally well simply by thinking over the problems (i.e., zero-shot).
@@
4.4 Reasoning-Related Results
@@
We observe that although LLMs can be competent zero-shot multi-problem solvers for reasoning, their performance becomes consistently worse than expected under multiple mixed-source problems. Similar to our arguments in last section, the consistent performance declines even with a small number (e.g., 2 or 4) of mixed-source problems may indicate a lack of human-like understanding, as LLMs’ reasoning capabilities are easily impacted by the mixing of problems from different sources.
@@
Figure 3: Average accuracy of GPT-3.5 and Llama-3 70B on multiple single-source (MultiReasonSS) and mixed-source (MultiReasonMS) reasoning problems. (A) MultiReasonSS results on 6 selected benchmarks for space reasons. The other 6 benchmarks show similar results (see Appendix C). (B) MultiReasonMS results across 6 distinct benchmark combinations, where each benchmark contributes a problem to each mixed-source prompt.
@@
5 Further Analyses
This section provide further analyses to better understand LLMs’ multiple problem handling capabilities, their limitation, and what enables such capabilities. It starts with two error analyses for BatchClf, aiming to see whether models make similar prediction errors and positional errors under BatchClf compared to SingleClf. We then investigate the reason why SelectAll appears to be harder than SelectOne, which seems counter-intuitive. Lastly, we explore model-level factors that may enable LLMs to receive and handle multiple problems at once.
5.1 BatchClf Error Analysis
Given the strong BatchClf results, two natural questions arise: do LLMs make similar errors under MPP and how do their errors distribute across positions? For each one of the 180 LLM/benchmark/task size combinations, we use chi-squared tests to compare the proportional error distribution across class labels under SingeClf and BatchClf and compute the cumulative error density across positions under BatchClf. We observe that (1) only in 9 out of 180 (or 5%) cases, error labels are distributed significantly differently between SingleClf and BatchClf (p < 0.05); and (2) surprisingly, LLMs typically do not display a clear positional bias or a serial position effect as known in psychology Murdock (1962), when solving sufficiently many problems at once (Fig 4). This is in contrast to previous studies based on single-problem prompts where LLMs are found to be better at using information from the beginning (primacy bias) or the end (receny bias) of the prompt Liu et al. (2024); Levy et al. (2024).
5.2 Why is SelectAll Harder than SelectOne
We investigate the reasons in Fig 5, which shows that when asked to select text indices for one class label at a time independently, LLMs almost always assign an index to multiple labels (i.e., contradiction) and leave some indices unselected (i.e., non-excluded middles) more often. This showcases a lack of internalized planning with modern zero-shot LLMs, although different LLMs may make these two types of errors in different proportions. In contrast, when LLMs have to select indices for all labels at once, they are less likely to generate directly illogical answers in a single output as their answer to the $(i+1)_{th}$ label is conditioned by their answer to the $i_{th}$ label.
5.3 Exploring Model-level Factors that may Enable MPP
Since so far we have only tested decoder-only and instruction-tuned LLMs, which all show strong performance under MPP, we explore if these two factors enable MPP. For these reasons, we test with greedy decoding (1) three FLAN-T5 models Chung et al. (2022), i.e., Large, XL, XXL; and (2) three pretrained LLMs, i.e., Llama-3 8B (Base, Meta, 2024), GPT-3 1.3B, and GPT-3 175B Brown et al. (2020). We run these 6 LLMs on CoLA at task sizes 1 and 5 under zero-shot settings with results shown in Table 3. Experiments with Coin Flips show similar results, but the LLM outputs are less meaningful, as discussed in Appendix D.2. We make the following observations from Table 3.
Instruction tuning helps.
This is because pretrained decoder-only LLMs either cannot handle multiple problems at once or their performance is much worse than their instruction-tuned counterparts. However, unlike what Laskar et al. (2023) suggests, instruction tuning may not be a necessary condition for MPP, since both Llama-3 8B and GPT-3 175B can perform reasonably well in BatchClf on CoLA.
FLAN-T5 can barely respond to MPP, regardless of model sizes.
We suspect that this may not be due to their encoder-decoder structures, but other factors such as training data and reinforcement learning from human feedback Christiano et al. (2017), which FLAN-T5 models lack. We leave it for future investigation.
Scaling model size seems helpful.
With other factors being identical, larger models appear to perform better than the smaller ones under MPP. For example, FLAN-T5-Large is outperformed by both FLAN-T5-XL and -XXL. Furthermore, while GPT-3 1.3B and FLAN-T5-XL can perform SingleClf close to or even better than GPT-3 175B and FLAN-T5-XXL, only the larger models can do BatchClf to varying extents–the two smaller models cannot do the task at all.
Final remark.
@@ 5.3 Exploring Model-level Factors that may Enable MPP
Table 3: SingleClf and BatchClf (task size 5) accuracy (%) of three pretrained LLMs and three FLAN-T5 models on CoLA. We also include the results of Llama-3 8B (Instruct) and GPT-3.5 (likely base model: GPT-3 175B) from Section 4.3 to compare with their respective base models. The last column is the average number of LLM-generated answers for BatchClf (expects 5). When it is 1, accuracy is not calculated to avoid overinterpretation.
@@ 6 Conclusion
In this study, we present a comprehensive and systematic MPE of LLMs. We evaluate various LLMs from 4 model families on single-source multi-problem prompts constructed from 6 classification and 12 reasoning benchmarks. In line with previous few-shot results, we confirm that LLMs are competent multi-problem solvers for classification and reasoning under zero-shot settings. Moreover, we find multiple pieces of evidence that validate the strong innate multiple problem handling capabilities of LLMs, such as the similar classification errors LLMs make under SPP and MPP, the lack of obvious positional biases, and the transferrability of zero-shot-CoT under MPP. Leveraging the strong multiple problem handling capabilities, we show that zero-shot MPP can be cost-efficient.
@@
Overall, our experiment demonstrate surprisingly consistent observations across different LLMs and across multi-problem prompts constructed from various benchmarks. This consistency indicates the reliability and fruitfulness of MPE as an evaluation paradigm.
As a result of our study, we create a new benchmark comprising 53,100 zero-shot multi-problem prompts. We call it ZeMPE, which stands for Zero-shot Multi-Problem Evaluation. We release ZeMPE to aid future MPE research.
Acknowledgements
We are grateful for the supports from the Institute for Advanced Computational Science (IACS) at Stony Brook University, in particular the free GPT access it provides. Zhengxiang Wang is supported by IACS’s Junior Researcher Award since Fall 2024. We thank three anonymous reviewers from the Generation, Evaluation & Metrics (GEM) Workshop 2025 for their helpful comments. The paper was presented at All Things Language And Computation (ATLAC) organized by the NLP reading group at Stony Brook University. We also thank the audience there for their discussions and feedback about the paper.
Zhengxiang Wang and Owen Rambow were supported in part by funding from the Defense Advanced Research Projects Agency (DARPA) under Contracts No. HR01121C0186, No. HR001120C0037, and PR No. HR0011154158. Any opinions, findings and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of DARPA.
@@ Appendix A LLM Details
We use a total of 13 LLMs in our study. Table 4 describes the specific versions for these LLMs and highlights their differences in terms of architecture, open weights, Supervised Fine-Tuning (SFT, Wei et al., 2022), and Reinforment Learning from Human Feedback (RLHF, Christiano et al., 2017) etc.
@@
Appendix B Classification-Related Results
This section contains additional details for Section 4.3. The prompt details for the experiments can be found in Appendix E.
@@ B.1 Full Results
The full results obtained from Section 4.3 are visualized in Fig 7. We exclude the results of Vicuna on AGNews when task size is 100 because the prompts exceed the model’s context length.
@@ B.3 Zero-shot MPP can be cost-efficient
Appendix C Reasoning-Related Results
This section provides additional details for Section 4.4. The prompt details for the experiments can be found in Appendix E.
@@ C.2 More Single-Source Results
Appendix D Further Analyses
D.1 Positional Errors under BatchClf
D.2 Exploring Model-level Factors that may Enable MPP
This section describes the results of the three pretrained base LLMs and three FLAN-T5 models on Coin Flips at task sizes 1 and 2 from Section 5.3, shown in Table 6. Similar to Table 3, Table 6 shows that instruction-tuned models perform much better on multi-problem prompts and that FLAN-T5 models can barely handle multi-problem prompts.